OpenStudy (anonymous):

A father said to his son, "I was as old as you are at the present at the time of your birth". If the father's age is 38 years now, the son's age five years back was????

OpenStudy (anonymous):

For these questions the tricky part is to find the equation that needs solving. In order to find it first you will need to decide what are you looking for. In other words what is x?

OpenStudy (anonymous):

Any idea? How should we define x?

OpenStudy (anonymous):

present age of son?

OpenStudy (anonymous):

Can be. Will you try to write and equation with it

OpenStudy (anonymous):

no:(

OpenStudy (anonymous):

I was as old as you are at the present at the time of your birth This needs to be translated to maths

OpenStudy (anonymous):

yes and this is the main problem.....actually i know the ans...but not the logic:(

OpenStudy (anonymous):

OK no problem. So the father is 38 and the son is x years old. How old was the father when his son was born? Use 38 and x

OpenStudy (anonymous):

2x=38 but how???i dont know

OpenStudy (anonymous):

First just figure this out: How old was the father when his son was born?

OpenStudy (anonymous):

I will give you a hint: my dad is 62 and I am 30. When I was born my dad was 32. How did I calculate this?

OpenStudy (anonymous):

19 years....but that's the point where i get confused....how can we find this by que that fathers age is double than that of his son..

OpenStudy (anonymous):

Can you answer my last question?

OpenStudy (anonymous):

ok let me think about it

OpenStudy (anonymous):

Sorry rosy I have to go for 5 minutes. Will be back later

OpenStudy (anonymous):

adding 32 in your age

OpenStudy (anonymous):

ok fine

OpenStudy (phi):

my dad is 62 and I am 30. *** adding 32 in your age *** yes because 30 + 32 = 62 now if we add -30 to both sides , we could also say 32 = 62 - 30 in other words, 30 years ago, the father was 32.

OpenStudy (phi):

father's age minus son's age is the age of the father when the son was born.

OpenStudy (anonymous):

I am back and phi is showing you the correct logic

OpenStudy (phi):

now notice we can say the same thing with letters my dad is 62 and I am x dad's age when son born is 62 - x

OpenStudy (anonymous):

but in my que 38-x=?

OpenStudy (phi):

so now try to figure out father's age when father's age is 38 years now son's age now is x years old how old was the father when the son was born ?

OpenStudy (phi):

38-x is the age of the father when the son was born.

OpenStudy (phi):

I was as old as you are now in other words, father's age (38-x) is the same as the son's age now (x) they are equal.

OpenStudy (anonymous):

and how can we find the value of x?

OpenStudy (anonymous):

oh ok...thank u:)

OpenStudy (phi):

can you write the equation?

OpenStudy (phi):

you should get 38-x = x

OpenStudy (anonymous):

but 62-30 not equal to 30 in andras exmp

OpenStudy (anonymous):

again confused:(

OpenStudy (phi):

andras uses different numbers. The thing that is the same is this: age of father minus age of son now equals age of father when the son was born.

OpenStudy (phi):

it turns out andras is not the same age as his father's age when andras was born. but in your problem, they tell us: son's age now ( that is, x) is the same as the father's age back then (38 - x)

OpenStudy (phi):

to get "unconfused" , convince yourself that if the father is 38 now, x years ago, he was 38-x (years old)

OpenStudy (phi):

if the father is 38 now, x years ago, he was 38-x (years old) does that part make sense?

OpenStudy (anonymous):

thaks a lot to both andras and phi....now it is clear to me

OpenStudy (phi):

once you get to 38-x= x you solve by add +x to both sides 38-x+x = x+x -x + x on the left side is 0. on the right side, x+x is 2x 38= 2x now divide both sides by 2 to get 19= x in other words, the son is 19 years old. the father is 38 years old. 19 years ago (when the son was born), the father was 38-19= 19 we see the father was 19 years old when his son was born... the same age as his son is *now*

OpenStudy (anonymous):

Great :)